The following information has been collected from Ms. Divya the head of personnel

department of ART Limited for the year ended 31st March 2020


Number of workers at the beginning of the year -8000

Number of workers at the end of the year -9600

Number of workers left the company during the year -500

Number of workers discharged during the year -100

Number of workers replaced due to leaving and discharge -700

Additional workers employed for expansion during the year -1500


You are required to calculate the labour turnover rate by using separation method,

replacement method and flux method, further discuss each of them briefly.


1
Expert's answer
2021-05-24T12:31:40-0400

Separation method

Labour Turnover = (separated workers during the period/ Average workers during the period ) x 100 %

Where: Average Workers = (workers at the beginning + workers at the end)/2

                                          (8000+9600)/2 = 8800

Therefore: 600/8800 X 100% =6 .818% hence this is the percentage number of workers left during the period.

Replacement method

Labour Turnover = number of workers replaced during the period/Average number of workers during the period

                           = (700/8800) X 100 % = 7.955% hence the percentage number of replaced workers during the year.

Flux method

Labour Turnover = (number of workers separated + number of workers replaced)/ Average number of workers during the year.

                           = 600 + 700 = 1300

                           = 1300/8800 X 100%

                            = 14.77 % hence the percentage combination of both separated and replaced workers during the year.

 

 

 

System

22:49

Your answer received. The quality department might check it. If there are any concerns, we will let you know.

22.05.21

You

19:29

Separation method

Labour Turnover = (separated workers during the period/ Average workers during the period ) x 100 %

Where: Average Workers = (workers at the beginning + workers at the end)/2

                                          (8000+9600)/2 = 8800

Therefore: 600/8800 X 100% = 6.818% hence this is the percentage number of workers left during the period.

Replacement method

Labour Turnover = number of workers replaced during the period/Average number of workers during the period

                           = (700/8800) X 100 % = 7.955% hence the percentage number of replaced workers during the year.

Flux method

Labour Turnover = (number of workers separated + number of workers replaced)/ Average number of workers during the year.

                           = 600 + 700 = 1300

                           = 1300/8800 X 100%

                            = 14.77 % hence the percentage combination of both separated and replaced workers during the year.